OpenStudy (anonymous):

What is the value of x? -2x + 8 = -16

OpenStudy (anonymous):

move 8 to the otherside, and divide by -2

OpenStudy (anonymous):

your answer is -12 I am positive

OpenStudy (anonymous):

No, I got 32 ..

OpenStudy (anonymous):

32 is not the answer its -12 do it again

OpenStudy (anonymous):

I am really GOOD AT MATH I am right

OpenStudy (anonymous):

No. Look. \[-2\times -12=24+8=32\]

OpenStudy (anonymous):

It's positive 12. G'day.

OpenStudy (anonymous):

yes sorry its 12positive now I am right SORRY

OpenStudy (anonymous):

-2x=-16-8 -2x=-24 -2x/2=-24/-2 x=12